OpenStudy (kitkat16):

Contestants in a game show are each given three pairs of numbers (for example, 2-3, 5-6, and 6-1). A person in the audience is randomly selected to roll two dice, one white and one red, where white symbolizes the first number in the pair and red the second number. What is the probability that each player has of winning?

OpenStudy (kitkat16):

Is it 2/3 ? anyone

OpenStudy (kitkat16):

@Preetha can you help me please?

OpenStudy (kitkat16):

@ganeshie8 can u help?

OpenStudy (kitkat16):

6 numbers with 12 possible outcomes yes?

ganeshie8 (ganeshie8):

is that the complete question ? whats the criteria for winning ?

OpenStudy (kitkat16):

but only 6 possible outcomes per dice

OpenStudy (kitkat16):

yes that is the complete question

OpenStudy (kitkat16):

my choices are 1/2 1/3 2/3 1/12 1/5 but Im confused how these apply

ganeshie8 (ganeshie8):

rolling two dies give you 6^2 = 36 total outcomes

ganeshie8 (ganeshie8):

out of that, only 3 are in your favor so probability = 3/36 = 1/12

OpenStudy (kitkat16):

I thought that but being that the 2 dice are different colors they are not combined?

OpenStudy (kitkat16):

white one have 6 octcomes and the red one having 6 outcomes

ganeshie8 (ganeshie8):

you will get the same answer by analyzing like that

ganeshie8 (ganeshie8):

the total outcomes are : white : 1 red : 1,2,3,4,5,6 ... white : 6 red : 1,2,3,4,5,6

ganeshie8 (ganeshie8):

that gives u 36 total outcomes ^

OpenStudy (kitkat16):

ok thanks im just over thinking it. Not good at these word problems

ganeshie8 (ganeshie8):

overthinking (any kind of thinking) is always good for these problems :)

OpenStudy (kitkat16):

lol thanks again :)

ganeshie8 (ganeshie8):

lets do this using ur idea : 6 total outcomes for white die, 3 are in favor so the probability for getting correct on white is 1/2

ganeshie8 (ganeshie8):

So you must have gotten one of the first numbers : 2-3, 5-6, and 6-1 on first die

OpenStudy (kitkat16):

yes

ganeshie8 (ganeshie8):

After that, there is only one favorable outcome for winning - the correct pair !

ganeshie8 (ganeshie8):

so probability for getting the matching pair on red die is : 1/6

ganeshie8 (ganeshie8):

the combined probability is : 1/2*1/6 = 1/12

ganeshie8 (ganeshie8):

working it either way is okay...

OpenStudy (kitkat16):

ok thanks the other way is easier though :)

ganeshie8 (ganeshie8):

yw

ganeshie8 (ganeshie8):

and yes the first method is shorter, but its a good to look these problems from different angles